To grow 100 kilograms of strawberries requires inputs that cost 80 euro. the same value of inputs would be needed to grow 1000 kilogram of wheat?
Answer #1

According to the question, the cost of growing 100 kilograms of strawberries is same as the cost of growing 1000 kilograms of wheat, that is, 80 euro. The average cost of production of strawberries is 80/100 = 0.8 euro per kilogram and the average cost of producing wheat is 80/1000 = 0.08 euro per kilogram. Thus the average cost of production of strawberries is 0.8/ 0.08 = 10 times the average cost of production of wheat.